PART 1: Glen Glutenfree is 24 years old, 5’6”, and weighs 198 pounds.  He wants to reduce...

PART 1:

Glen Glutenfree is 24 years old, 5’6”, and weighs 198 pounds.  He wants to reduce his weight to the recommended BMI range.  Currently, he walks for 40 minutes three days per week.  He also lifts weights 2 days a week for about 20 minutes per session.

A.  How much weight does he need to lose to reach the recommended minimum BMI?

B.  Using the Mifflin St. Jeor Equation worksheet below, calculate the number of calories he needs to meet his goal weight:     

Step 1:  Calculate calorie needs for goal weight

Equation for Males

(4.55 x wt lbs) + (15.88 x ht in) – (5 x age in years) + 5 = __________calories

Equation for Females

(4.55 x wt lbs) + (15.88 x ht in) – (5 x age in years) – 161 = ___________calories

Step 2: Multiply the number from Step 1 by the appropriate Activity Factor (see activity factors below):  ___________________calories

1.1 = Sedentary = routine daily activities only

1.2 = Light = at least 30 min of intentional lifestyle activity or exercise 3 to 5 days per week

1.3 = Moderate = at least 30 min of moderate-intensity activity or exercise, such as brisk walking, aerobic dance,

or yard work, on 5 or more days of the week

1.4 = Vigorous = at least 30 minutes vigorous-intensity activity, such as running or fast bicycling, on 5 or more

days of the week

1.55 (Very Vigorous Activity) – greater than 30 minutesof vigorous intensity activity, such as running      or fast bicycling on 5 or more days per week

PART 2:  

What percentage of his recommended calorie level should come fromprotein?   ______% to ______ %

PART 3:

Calculate the number of calories this represents from PART 2:

______________ to ___________ calories

PART 4:

How many grams of protein does this represent?

_______ to ______ grams of protein

PART 5:

Does his current level of physical activity and exercise meet the minimum requirements for health benefits?  

PART 6:

Do you have any specific suggestions that could increase Mr. Glutenfree’s chance of reaching his goals.
Answer #1

Age of Glen gluten-free is 24years

Height of the patient is 5feet 6inches= 1.67m

Weight of the patient is 198lbs = 89.1kg

BMI = WEIGHT IN KG /HEIGHT IN METER SQUARE

= 89.1 / 1.67× 1.67

= 32kg /meter square

Normal BMI is 18.5 to 24.9

Step 1.

A) The amount of weight he should lose for minimum BMI is around 33kg because for 5feet 6inch height the normal weight is 53 kg . So, he should lose 33kgs.

B) Calories needed for goal weight for males :-

Equation is (4.55× wt. in lbs) +( 15.88×ht. in inches) - ( 5× age in years ) +5

=( 4.55× 198)+(15.88× 66) -(5×24) +5

= (900.9) +(1048) -(120) +5

= (1948.9) -(120) +5

= 1823.9 calories

Equation for females is (4.55×wt.in lbs ) +(15.88××ht in inches) -(5× age in years )-161

= (4.55×198) +(15.88×66) -(5×24)-161

= (900.9)+(1048)-(120)- 161

= 1667.9 calories

Step 2. - multiply the calories from appropriate activity factor

The appropriate activity factor is 1.2 for Glen gluten-free because he walks for 40mins three days a week and weight lifting two days a week for 20mins

Calories of Glen gluten-free is 1823.9 calories

Multiply it by 1.2 = 1823.9×1.2

= 2188.68 calories

# Part 2:

1gm of protein provides 4kcal

Total recommended calories of Glen gluten-free is 1823.9 calories

Then protein will be 1823.9/4

= 455.9kcal

= 455.9/1823.9 ×100

= 24.9 % of protein

= 24 %to 24.9%

#Part 3:

= The number of calories from protein

= 1823.9/4

= 455 to 456 kcal

Part 4:

Grams of protein :-

4kcal is provided by 1gm of protein

1kcal is provided by 1/4 gm

Then 455.9 kcal is provided by 1/4 × 455.9

= 113.9 gm of protein

Range of protein in gm is 113 to 115 gm